One base of a trapezoid is 12 in. long and the altitude is 8 in. long. If the area is 88 sq. in., find the length of the other base.
Answer by malakumar_kos@yahoo.com(315)


one base of the trapezoid a=12 in
the other base b=x in
altitude h=8.in
Area A=88 sq.in
A = 1/2.h.(a+b) 88=1/2.8.(12+x) 88/4 = 12+x 22=x+12 x=10 in
the other base of the trapezoid = 10.in